Circular No. 6341

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 E1 E. F. Helin and K. J. Lawrence, Jet Propulsion Laboratory, on behalf

of the Near-Earth-Asteroid Tracking (NEAT) Team (E. Helin, S. Pravdo, S. Groom, C. Clark, R. Bambery, S. Levin, J. Lorre, S. Shaklan, and K. Lawrence), report the discovery of a comet on images obtained with the 1-m f/2.2 GEODSS telescope at Haleakala by the on-site team (R. Byrd, A. Esquibel, C. Cotton, and D. Bascon). The comet is diffuse with strong condensation and a tail 10′′–15′′ long. Available CCD observations:

1996 UT  α_{2000}  δ_{2000}  m_{1} Observer
Mar. 15.25450 9^{h}01^{m}13.^{s}54 +18^{o}58′39′′.3 16.2 Helin
15.27535 9 01 10.04 +18 59 26.6 16.3  _{”}
15.29746 9 01 06.28 +19 00 17.2 16.3  _{”}
16.16319 8 58 42.63 +19 32 55.1 16.0 Rogers
16.17153 8 58 41.37 +19 33 11.9 16.0  _{”}
16.18752 8 58 38.69 +19 33 50.0 16.1  _{”}
16.21910 8 58 33.36 +19 35 01.3 16.1  _{”}

E. F. Helin and K. J. Lawrence (Haleakala-NEAT/GEODSS). J. E. Rogers, E. F. Helin, R. P. Helin, B. G. Marsden, J. B. Marsden, S. N.

Marsden, G. Zappa, and D. Zappa (Camarillo). 0.25-m reflector.

COMET C/1996 B2 (HYAKUTAKE) R. M. West, European Southern Observatory (ESO), writes: “Two

exposures (26 and 60 min; Kodak 4415 film + GG385 filter), obtained on Mar. 14.3 UT by G. Pizarro with the 1-m ESO Schmidt telescope, show a further development of the tail system; it is now morphologically similar to that seen in other major comets at the present heliocentric distance and magnitude. There are at least ten straight streamers in the sector p.a. 263^{o}– 310^{o}, some reaching to 75′. The main ion tail at p.a. 285^{o} has numerous kinks (e.g., at 75′ and 108′ distance) and bifurcations; its length exceeds 300′ (field border). There is a fainter, diffuse and broad (supposedly dust) tail underlying the ion tail and also longer than 300′; at this distance, its total width is more than 100′.”

Visual m_{1} estimates by J. E. Bortle, Stormville, NY: Mar. 12.33 UT, 4.3 (10×50 binoculars); 13.26, 4.2 (10×50 binoculars); 16.28, 2.6 (naked eye; coma diameter 0^{o}.9).

1996 March 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green